OpenStudy (anonymous):

14 is 28% of what number

OpenStudy (texaschic101):

percentage formula : is/of = %/100 14 is 28% of what number. is = 14 % = 28 of = x now sub is/of = %/100 14/x = 28/100 (cross multiply since this is a proportion) (28)(x) = (14)(100) 28x = 1400 x = 1400/28 x = 50 check.. 14 is 28% of 50 14 = 0.28(50) 14 = 14 (correct) so your answer is 50 you can always use this formula for percentage problems, just be sure and sub the numbers in the right place.